If 84players split themselves into teams, how many more teams can they form by splitting into teams if 4 instead of teams of 6.
A. 5
B. 6
C. 7
D, 14

Answer by jim_thompson5910(35256)

Teams of 4:
84/4 = 21
so we have 21 teams of four

Teams of 6:
84/6 = 14
we have 14 teams of six


Difference:
21-14 = 7

So we have 7 more teams if we go with teams of four, instead of teams of six.

Answer: C) 7